Could someone help me with this problem? Thanks!
Solve
9x^4-18x^2+5=0
The solution is x=[ ]

Answer by scott8148(6628)

factoring ___ (3x^2 - 5)(3x^2 - 1) = 0

3x^2 - 5 = 0 ___ x = ± sqrt(5/3)

3x^2 - 1 = 0 ___ x = ± sqrt(1/3)